\begin{thebibliography}{52}
\expandafter\ifx\csname natexlab\endcsname\relax\def\natexlab#1{#1}\fi

\bibitem[{Absil et~al.(2008)Absil, Mahony, and Sepulchre}]{absil08a}
Pierre-Antoine Absil, Robert Mahony, and Rodolphe Sepulchre. 2008.
\newblock \emph{Optimization Algorithms on Matrix Manifolds}.
\newblock Princeton University Press, Princeton, NJ.

\bibitem[{Adel et~al.(2017)Adel, Zhao, and Wong}]{adel17a}
Tameem Adel, Han Zhao, and Alexander Wong. 2017.
\newblock Unsupervised domain adaptation with a relaxed covariate shift
  assumption.
\newblock In \emph{Proceedings of the AAAI Conference on Artificial
  Intelligence}.

\bibitem[{Alaux et~al.(2019)Alaux, Grave, Cuturi, and Joulin}]{alaux19a}
Jean Alaux, Edouard Grave, Marco Cuturi, and Armand Joulin. 2019.
\newblock Unsupervised hyperalignment for multilingual word embeddings.
\newblock In \emph{Proceedings of the International Conference on Learning
  Representations}.
\newblock {URL:
  \url{https://github.com/facebookresearch/fastText/tree/master/alignment}}.

\bibitem[{Alvarez-Melis and Jaakkola(2018)}]{alvares18a}
David Alvarez-Melis and Tommi~S. Jaakkola. 2018.
\newblock Gromov-wasserstein alignment of word embedding spaces.
\newblock In \emph{Proceedings of the Conference on Empirical Methods in
  Natural Language Processing}.
\newblock {URL: https://github.com/dmelis/otalign}.

\bibitem[{Artetxe et~al.(2016)Artetxe, Labaka, and Agirre}]{artetxe16a}
Mikel Artetxe, Gorka Labaka, and Eneko Agirre. 2016.
\newblock Learning principled bilingual mappings of word embeddings while
  preserving monolingual invariance.
\newblock In \emph{Proceedings of the Conference on Empirical Methods in
  Natural Language Processing}, pages 2289--2294.

\bibitem[{Artetxe et~al.(2018{\natexlab{a}})Artetxe, Labaka, and
  Agirre}]{artetxe18a}
Mikel Artetxe, Gorka Labaka, and Eneko Agirre. 2018{\natexlab{a}}.
\newblock Generalizing and improving bilingual word embedding mappings with a
  multi-step framework of linear transformations.
\newblock In \emph{Proceedings of the AAAI Conference on Artificial
  Intelligence}, pages 5012--5019.

\bibitem[{Artetxe et~al.(2018{\natexlab{b}})Artetxe, Labaka, and
  Agirre}]{artetxe18b}
Mikel Artetxe, Gorka Labaka, and Eneko Agirre. 2018{\natexlab{b}}.
\newblock A robust self-learning method for fully unsupervised cross-lingual
  mappings of word embeddings.
\newblock In \emph{Proceedings of the Annual Meeting of the Association for
  Computational Linguistics}, pages 789--798.
\newblock {URL: https://github.com/artetxem/vecmap}.

\bibitem[{Artetxe et~al.(2018{\natexlab{c}})Artetxe, Labaka, and
  Agirre}]{artetxe18c}
Mikel Artetxe, Gorka Labaka, and Eneko Agirre. 2018{\natexlab{c}}.
\newblock Unsupervised statistical machine translation.
\newblock In \emph{Proceedings of the 2018 Conference on Empirical Methods in
  Natural Language Processing}, pages 3632--3642.

\bibitem[{Baktashmotlagh et~al.(2013)Baktashmotlagh, Harandi, Lovell, and
  Salzmann}]{baktashmotlagh13a}
Mahsa Baktashmotlagh, Mehrtash~T Harandi, Brian~C Lovell, and Mathieu Salzmann.
  2013.
\newblock Unsupervised domain adaptation by domain invariant projection.
\newblock In \emph{Proceedings of the IEEE International Conference on Computer
  Vision}, pages 769--776.

\bibitem[{Ben-David et~al.(2007)Ben-David, Blitzer, Crammer, and
  Pereira}]{ben07a}
Shai Ben-David, John Blitzer, Koby Crammer, and Fernando Pereira. 2007.
\newblock Analysis of representations for domain adaptation.
\newblock In \emph{Advances in neural information processing systems}, pages
  137--144.

\bibitem[{Bergmann(2019)}]{bergmann19a}
Ronny Bergmann. 2019.
\newblock {Optimisation on Manifolds in Julia}.
\newblock \url{https://github.com/kellertuer/Manopt.jl}.

\bibitem[{Bojanowski et~al.(2017)Bojanowski, Grave, Joulin, and
  Mikolov}]{bojanowski17}
Piotr Bojanowski, Edouard Grave, Armand Joulin, and Tomas Mikolov. 2017.
\newblock Enriching word vectors with subword information.
\newblock \emph{Transactions of the Association for Computational Linguistics},
  5:135--146.

\bibitem[{Borgwardt et~al.(2006)Borgwardt, Gretton, Rasch, Kriegel,
  Sch{\"o}lkopf, and Smola}]{borgwardt06a}
Karsten~M Borgwardt, Arthur Gretton, Malte~J Rasch, Hans-Peter Kriegel,
  Bernhard Sch{\"o}lkopf, and Alex~J Smola. 2006.
\newblock Integrating structured biological data by kernel maximum mean
  discrepancy.
\newblock \emph{Bioinformatics}, 22(14):e49--e57.

\bibitem[{Boumal et~al.(2014)Boumal, Mishra, Absil, and Sepulchre}]{boumal14a}
Nicolas Boumal, Bamdev Mishra, Pierre-Antoine Absil, and Rodolphe Sepulchre.
  2014.
\newblock Manopt, a {M}atlab toolbox for optimization on manifolds.
\newblock \emph{Journal of Machine Learning Research}, 15(Apr):1455--1459.

\bibitem[{Conneau et~al.(2018)Conneau, Lample, Ranzato, Denoyer, and
  J{\'e}gou}]{conneau18a}
Alexis Conneau, Guillaume Lample, Marc'Aurelio Ranzato, Ludovic Denoyer, and
  Herv{\'e} J{\'e}gou. 2018.
\newblock Word translation without parallel data.
\newblock In \emph{Proceedings of the International Conference on Learning
  Representations}.
\newblock {URL: https://github.com/facebookresearch/MUSE}.

\bibitem[{Cuturi(2013)}]{cuturi13a}
Marco Cuturi. 2013.
\newblock Sinkhorn distances: Lightspeed computation of optimal transport.
\newblock In \emph{Advances in neural information processing systems}, pages
  2292--2300.

\bibitem[{Daum{\'e}~III(2007)}]{daumeiii07a}
Hal Daum{\'e}~III. 2007.
\newblock Frustratingly easy domain adaptation.
\newblock In \emph{Proceedings of the Annual Meeting of the Association of
  Computational Linguistics}, pages 256--263.

\bibitem[{Douik and Hassibi(2019)}]{douik19a}
Ahmed Douik and Babak Hassibi. 2019.
\newblock Manifold optimization over the set of doubly stochastic matrices: A
  second-order geometry.
\newblock \emph{IEEE Transactions on Signal Processing}, 67(22):5761--5774.

\bibitem[{Edelman et~al.(1998)Edelman, Arias, and Smith}]{edelman98a}
Alan Edelman, Tom\'as~A. Arias, and Steven~T. Smith. 1998.
\newblock The geometry of algorithms with orthogonality constraints.
\newblock \emph{SIAM Journal on Matrix Analysis and Applications},
  20(2):303--353.

\bibitem[{Faruqui and Dyer(2014)}]{faruqui14a}
Manaal Faruqui and Chris Dyer. 2014.
\newblock Improving vector space word representations using multilingual
  correlation.
\newblock In \emph{Proceedings of the Conference of the European Chapter of the
  Association for Computational Linguistics}, pages 462--471.

\bibitem[{Fukumizu et~al.(2007)Fukumizu, Bach, and Gretton}]{fukumizu07a}
Kenji Fukumizu, Francis~R Bach, and Arthur Gretton. 2007.
\newblock Statistical consistency of kernel canonical correlation analysis.
\newblock \emph{Journal of Machine Learning Research}, 8(Feb):361--383.

\bibitem[{{Gopalan} et~al.(2011){Gopalan}, {Ruonan Li}, and
  {Chellappa}}]{gopalan11a}
R.~{Gopalan}, {Ruonan Li}, and R.~{Chellappa}. 2011.
\newblock Domain adaptation for object recognition: An unsupervised approach.
\newblock In \emph{International Conference on Computer Vision}.

\bibitem[{Grave et~al.(2019)Grave, Joulin, and Berthet}]{grave18}
Edouard Grave, Armand Joulin, and Quentin Berthet. 2019.
\newblock Unsupervised alignment of embeddings with {Wasserstein Procrustes}.
\newblock In \emph{International Conference on Artificial Intelligence and
  Statistics}.

\bibitem[{Hoshen and Wolf(2018)}]{hoshen18}
Yedid Hoshen and Lior Wolf. 2018.
\newblock Non-adversarial unsupervised word translation.
\newblock Technical report, arXiv preprint arXiv:1801.06126v3.

\bibitem[{Huang et~al.(2016)Huang, Absil, Gallivan, and Hand}]{huang16a}
Wen Huang, Pierre-Antoine Absil, Kyle~A. Gallivan, and Paul Hand. 2016.
\newblock Roptlib: an object-oriented {C}++ library for optimization on
  {Riemannian} manifolds.
\newblock Technical Report FSU16-14.v2, Florida State University.

\bibitem[{Jawanpuria et~al.(2019)Jawanpuria, Balgovind, Kunchukuttan, and
  Mishra}]{mjaw19a}
Pratik Jawanpuria, Arjun Balgovind, Anoop Kunchukuttan, and Bamdev Mishra.
  2019.
\newblock Learning multilingual word embeddings in latent metric space: A
  geometric approach.
\newblock \emph{Transactions of the Association for Computational Linguistics},
  7:107--120.

\bibitem[{Jawanpuria et~al.(2020)Jawanpuria, Meghwanshi, and Mishra}]{mjaw20a}
Pratik Jawanpuria, Mayank Meghwanshi, and Bamdev Mishra. 2020.
\newblock A simple approach to learning unsupervised multilingual embeddings.
\newblock Technical report, arXiv preprint arXiv:2004.05991.

\bibitem[{Joulin et~al.(2018)Joulin, Bojanowski, Mikolov, Grave, and
  J{\'e}gou}]{joulin18b}
Armand Joulin, Piotr Bojanowski, Tomas Mikolov, Edouard Grave, and Herv{\'e}
  J{\'e}gou. 2018.
\newblock Loss in translation: Learning bilingual word mapping with a retrieval
  criterion.
\newblock In \emph{Proceedings of the Conference on Empirical Methods in
  Natural Language Processing}.

\bibitem[{Klementiev et~al.(2012)Klementiev, Titov, and
  Bhattarai}]{klementiev12a}
Alexandre Klementiev, Ivan Titov, and Binod Bhattarai. 2012.
\newblock Inducing crosslingual distributed representations of words.
\newblock In \emph{Proceedings of the International Conference on Computational
  Linguistics: Technical Papers}, pages 1459--1474.

\bibitem[{Lebanon and Lafferty(2004)}]{lebanon04a}
Guy Lebanon and John Lafferty. 2004.
\newblock Hyperplane margin classifiers on the multinomial manifold.
\newblock In \emph{Proceedings of the International Conference on Machine
  learning}, page~66.

\bibitem[{Mahadevan et~al.(2018)Mahadevan, Mishra, and Ghosh}]{mahadevan18a}
Sridhar Mahadevan, Bamdev Mishra, and Shalini Ghosh. 2018.
\newblock A unified framework for domain adaptation using metric learning on
  manifolds.
\newblock In \emph{Joint European Conference on Machine Learning and Knowledge
  Discovery in Databases}, pages 843--860.

\bibitem[{Meghwanshi et~al.(2018)Meghwanshi, Jawanpuria, Kunchukuttan, Kasai,
  and Mishra}]{Meghwanshi_arXiv_2018}
Mayank Meghwanshi, Pratik Jawanpuria, Anoop Kunchukuttan, Hiroyuki Kasai, and
  Bamdev Mishra. 2018.
\newblock Mctorch, a manifold optimization library for deep learning.
\newblock Technical report, arXiv preprint arXiv:1810.01811.

\bibitem[{M\'emoli(2011)}]{memoli11a}
Facundo M\'emoli. 2011.
\newblock {G}romov--{W}asserstein distances and the metric approach to object
  matching.
\newblock \emph{Foundations of computational mathematics}, 11(4):417--487.

\bibitem[{Mikolov et~al.(2013)Mikolov, Le, and Sutskever}]{mikolov13a}
Tomas Mikolov, Quoc~V Le, and Ilya Sutskever. 2013.
\newblock Exploiting similarities among languages for machine translation.
\newblock Technical report, arXiv preprint arXiv:1309.4168.

\bibitem[{Peyr\'e and Cuturi(2019)}]{peyre19a}
Gabriel Peyr\'e and Marco Cuturi. 2019.
\newblock Computational optimal transport.
\newblock \emph{Foundations and Trends in Machine Learning}, 11(5-6):355--602.

\bibitem[{Peyr\'e et~al.(2016)Peyr\'e, Cuturi, and Solomon}]{peyre16a}
Gabriel Peyr\'e, Marco Cuturi, and Justin Solomon. 2016.
\newblock {G}romov-{W}asserstein averaging of kernel and distance matrices.
\newblock In \emph{Proceedings of the International Conference on Machine
  Learning}.

\bibitem[{Prettenhofer and Stein(2011)}]{prettenhofer11a}
Peter Prettenhofer and Benno Stein. 2011.
\newblock Cross-lingual adaptation using structural correspondence learning.
\newblock \emph{ACM Transactions on Intelligent Systems and Technology},
  3(1):13.

\bibitem[{Ruder(2019)}]{ruder19a}
Sebastian Ruder. 2019.
\newblock \emph{Neural transfer learning for natural language processing}.
\newblock Ph.D. thesis, National University of Ireland, Ireland.

\bibitem[{Sch{\"o}nemann(1966)}]{schonemann1966a}
Peter~H Sch{\"o}nemann. 1966.
\newblock A generalized solution of the orthogonal {Procrustes} problem.
\newblock \emph{Psychometrika}, 31(1):1--10.

\bibitem[{Smith(1994)}]{smith94a}
S.~T. Smith. 1994.
\newblock Optimization techniques on {R}iemannian manifold.
\newblock In A.~Bloch, editor, \emph{Hamiltonian and Gradient Flows, Algorithms
  and Control}, volume~3, pages 113--136. American Mathematical Society,
  Providence, RI.

\bibitem[{Smith et~al.(2017)Smith, Turban, Hamblin, and Hammerla}]{smith17a}
Samuel~L. Smith, David H.~P. Turban, Steven Hamblin, and Nils~Y. Hammerla.
  2017.
\newblock Offline bilingual word vectors, orthogonal transformations and the
  inverted softmax.
\newblock In \emph{Proceedings of the International Conference on Learning
  Representations}.

\bibitem[{Sun et~al.(2016)Sun, Feng, and Saenko}]{sun16a}
Baochen Sun, Jiashi Feng, and Kate Saenko. 2016.
\newblock Return of frustratingly easy domain adaptation.
\newblock In \emph{Proceedings of the AAAI Conference on Artificial
  Intelligence}.

\bibitem[{Sun et~al.(2015)Sun, Gao, Hong, Mishra, and Yin}]{sun15a}
Yanfeng Sun, Junbin Gao, Xia Hong, Bamdev Mishra, and Baocai Yin. 2015.
\newblock Heterogeneous tensor decomposition for clustering via manifold
  optimization.
\newblock \emph{IEEE transactions on pattern analysis and machine
  intelligence}, 38(3):476--489.

\bibitem[{Townsend et~al.(2016)Townsend, Koep, and Weichwald}]{townsend16a}
James Townsend, Niklas Koep, and Sebastian Weichwald. 2016.
\newblock Pymanopt: A python toolbox for optimization on manifolds using
  automatic differentiation.
\newblock \emph{Journal of Machine Learning Research}, 17(137):1--5.
\newblock {URL: https://pymanopt.github.io}.

\bibitem[{Villani(2009)}]{villani09}
C\'{e}dric Villani. 2009.
\newblock \emph{Optimal Transport: Old and New}.
\newblock Springer-Verlag.

\bibitem[{Vuli\'{c} and Moens(2015)}]{vulic15a}
Ivan Vuli\'{c} and Marie-Francine Moens. 2015.
\newblock Monolingual and cross-lingual information retrieval models based on
  (bilingual) word embeddings.
\newblock In \emph{Proceedings of the ACM SIGIR Conference on Research and
  Development in Information Retrieval}, pages 363--372.

\bibitem[{Wan et~al.(2011)Wan, Pan, and Li}]{wan11a}
Chang Wan, Rong Pan, and Jiefei Li. 2011.
\newblock Bi-weighting domain adaptation for cross-language text
  classification.
\newblock In \emph{International Joint Conference on Artificial Intelligence}.

\bibitem[{Wang et~al.(2015)Wang, Arora, Livescu, and Srebro}]{wang15a}
Weiran Wang, Raman Arora, Karen Livescu, and Nathan Srebro. 2015.
\newblock Stochastic optimization for deep cca via nonlinear orthogonal
  iterations.
\newblock In \emph{Annual Allerton Conference on Communication, Control, and
  Computing}, pages 688--695.

\bibitem[{Xu et~al.(2018)Xu, Yang, Otani, and Wu}]{xu18a}
Ruochen Xu, Yiming Yang, Naoki Otani, and Yuexin Wu. 2018.
\newblock Unsupervised cross-lingual transfer of word embedding spaces.
\newblock In \emph{Proceedings of the Conference on Empirical Methods in
  Natural Language Processing}.

\bibitem[{Zhang et~al.(2017{\natexlab{a}})Zhang, Liu, Luan, and Sun}]{zhang17a}
Meng Zhang, Yang Liu, Huanbo Luan, and Maosong Sun. 2017{\natexlab{a}}.
\newblock Adversarial training for unsupervised bilingual lexicon induction.
\newblock In \emph{Proceedings of the Annual Meeting of the Association for
  Computational Linguistics}, pages 1959--1970.

\bibitem[{Zhang et~al.(2017{\natexlab{b}})Zhang, Liu, Luan, and Sun}]{zhang17b}
Meng Zhang, Yang Liu, Huanbo Luan, and Maosong Sun. 2017{\natexlab{b}}.
\newblock Earth mover's distance minimization for unsupervised bilingual
  lexicon induction.
\newblock In \emph{Proceedings of the Conference on Empirical Methods in
  Natural Language Processing}, pages 1934--1945.

\bibitem[{Zhou et~al.(2019)Zhou, Ma, Wang, and Neubig}]{zhou19a}
Chunting Zhou, Xuezhe Ma, Di~Wang, and Graham Neubig. 2019.
\newblock Density matching for bilingual word embedding.
\newblock In \emph{Proceedings of the Conference of the North American Chapter
  of the Association for Computational Linguistics}.

\end{thebibliography}

\documentclass[11pt,a4paper]{article}
\usepackage[hyperref]{acl2020} 
\usepackage{times}
\usepackage{latexsym}

\usepackage{microtype}

\aclfinalcopy 
\def\aclpaperid{2728} 

\usepackage{xspace,mfirstuc,tabulary}
\usepackage{breakurl}

\author{
 Pratik Jawanpuria, \ Mayank Meghwanshi, \ Bamdev Mishra\\
 Microsoft, India\\
 \texttt{\{pratik.jawanpuria,mamegh,bamdevm\}@microsoft.com}\\
}

\date{}

\usepackage{enumitem}
\setlist{nolistsep}
\usepackage{nicefrac}

\usepackage{booktabs}
\usepackage{mathtools}
\usepackage{color,multirow,wrapfig}
\usepackage[normalem]{ulem}

\usepackage{graphicx} 
\usepackage{epsfig} 
\usepackage{subfigure} 
\usepackage{caption}

\usepackage{algorithm}
\usepackage{algorithmic}

\usepackage{amssymb}
\usepackage{amsmath}
\usepackage{amsthm}
\usepackage{mathtools}

\newcommand{\thead}[1]{\multicolumn{1}{c}{\textbf{#1}}}
\newcommand{\lhead}[1]{\multicolumn{1}{l}{\textbf{#1}}} 

\newcommand{\algname}{\mbox{MBA}}

\newcommand{\mat}[1]{{\bf #1}}
\def\R{\mathbb{R}}
\newcommand{\norm}[1]{\left\|#1\right\|}

\def\bA{\mathbf A}

\def\bC{\mathbf C}
\def\bD{\mathbf D}
\def\bI{{\mathbf I}}

\def\bU{{\mathbf U}}
\def\bV{{\mathbf V}}
\def\bW{{\mathbf W}}
\def\bX{{\mathbf X}}
\def\bY{{\mathbf Y}}
\def\bZ{{\mathbf Z}}

\def\DS{{\mathbb{DS}}}

\def\minop{\mathop{\rm min}\limits}

\title{Geometry-aware Domain Adaptation for\\ Unsupervised Alignment of Word Embeddings}
\begin{document} 
\maketitle

\begin{abstract} 
We propose a novel manifold based geometric approach for learning unsupervised alignment of word embeddings between the source and the target languages. Our approach formulates the alignment learning problem as a domain adaptation problem over the manifold of doubly stochastic matrices. 
This viewpoint arises from the aim to align the second order information of the two language spaces. 
The rich geometry of the doubly stochastic manifold allows to employ efficient Riemannian conjugate gradient algorithm for the proposed formulation.
Empirically, the proposed approach outperforms state-of-the-art optimal transport based approach on the bilingual lexicon induction task across several language pairs. The performance improvement is more significant for distant language pairs.
\end{abstract}

\section{Introduction}
Learning bilingual word embeddings is an important problem in natural language processing  \citep{mikolov13a,faruqui14a,artetxe16a,conneau18a}, with usage in cross-lingual information retrieval \citep{vulic15a}, text classification \citep{wan11a,klementiev12a}, machine translation \citep{artetxe18c}  etc. 
Given a source-target language pair, the aim is to represent the words in both languages in a common embedding space. This is usually achieved by learning a linear function that maps word embeddings of one language to the embedding space of the other language \citep{mikolov13a}. 

Several works have focused on learning such bilingual mapping in supervised setting, using a bilingual dictionary during the training phase  \citep{artetxe18a,joulin18b,mjaw19a}. Recently, unsupervised bilingual word embeddings have also been explored  \citep{zhang17a,zhang17b,conneau18a,artetxe18b,hoshen18,grave18,alvares18a,zhou19a,mjaw20a}. 



Learning unsupervised cross-lingual mapping may be viewed as an instance of the more general unsupervised domain adaptation problem \citep{ben07a,gopalan11a,sun16a,mahadevan18a}. The latter fundamentally aims at aligning the input feature (embeddings) distributions of the source and target domains  (languages). 
In this paper, we take this point of view and learn cross-lingual word alignment by finding alignment between the second order statistics of the source and the target language embedding space. 

We formulate a novel optimization problem on the set of doubly stochastic matrices. The objective function consists of matching covariances of words from source to target languages in a least-squares sense. For optimization, we exploit the fact that the set of doubly stochastic matrices has rich geometry and forms a Riemannian manifold \citep{douik19a}. The Riemannian optimization framework \citep{absil08a,edelman98a,smith94a} allows to propose a computationally efficient conjugate gradient algorithm \citep{douik19a}. Experiments show the efficacy of the proposed approach on the bilingual lexicon induction benchmark, especially on the  language pairs involving distant languages. 


\section{Motivation and Related Work} \label{sec:related_work}
We introduce the bilingual word alignment setup followed by a discussion on domain adaptation approaches.

\noindent\textbf{Bilingual alignment.} Let $\bX\in\R^{n\times d}$ and $\bZ\in\R^{n\times d}$ be $d$-dimensional word embeddings of $n$ words of the source and the target languages, respectively. The aim is to learn a linear operator $\bW: \R^d\rightarrow\R^d$ that best approximates source embeddings in the target  language space. 

In the supervised setup, a list of source words and their translations in the target language is provided. This is represented by an \emph{alignment} matrix $\bY$ of size $n\times n$, where $\bY_{ij}=1$ if $j$-th word in the target language is a translation of the $i$-th word in the source language and $\bY_{ij}=0$ otherwise. 
A standard way to learn orthogonal $\bW$ is by solving the \textit{orthogonal Procrustes} problem \citep{artetxe16a,smith17a}, i.e., 
\begin{equation} \label{eq:procrustes}
\begin{array}{lll}
\minop_{\bW \in\R^{d\times d}} & \norm{\bX \bW - \bY \bZ}_{\rm Fro}^2\\
{\rm subject\ to} & \bW^\top \bW = \bI,
\end{array}
\end{equation}
where $\norm{\cdot}_{\rm Fro}$ is the Frobenius norm and $\bI$ is the identity matrix. Problem (\ref{eq:procrustes}) has the closed-form solution $\bW^\star = \bU\bV^\top$, where $\bU$ and $\bV$ are the respective left and right orthogonal factors of the singular value decomposition of $\bX^\top \bY \bZ$ \citep{schonemann1966a}. 

In the unsupervised setting, $\bY$ is additionally unknown apart from $\bW$. Most unsupervised works \citep{zhang17b,artetxe18b,grave18,conneau18a} tackle this challenge by learning $\bY$ and $\bW$ jointly. However, their performance rely on finding a good initialization candidate for the alignment matrix $\bY$ \cite{zhang17b,grave18,alaux19a,mjaw20a}. 



Performing optimization over the set of binary matrices, $\bY\in\{0,1\}^{n\times n}$, to learn the bilingual alignment matrix is computationally hard. Hence, some    works \citep{zhang17b,xu18a} view the source and the target word embedding spaces as two distributions and learn $\bY$ as the transformation that makes the two distributions close. 
This viewpoint is based on the theory of \emph{optimal transport} \citep{villani09,peyre19a}. $\bY$ is, thus, modeled as a \emph{doubly stochastic} matrix: the entries in $\bY \in [0,1]$ and each row/column sums to $1$. 
Permutation matrices are extreme points in the space of doubly stochastic matrices.

\citet{alvares18a} propose learning the doubly stochastic $\bY$ as a transport map between the \emph{metric spaces} of the words in the source and the target languages. 
They optimize the {Gromov-Wasserstein} (GW) distance, which measures how distances between pairs of words are mapped across languages.
For learning $\bY$, they propose to
\begin{equation}\label{eq:GW}
\begin{array}{lll}
\minop_{\bY \in \DS^n} & - {\rm Trace}(\bY^\top \bC_\bX \bY \bC_\bZ) ,
\end{array}
\end{equation}
where $\DS^n \coloneqq \{ \bY \in \R^{n\times n} : \bY \geq 0,  \bY^\top \mat{1} = \mat{1} \text{ and } \bY \mat{1} = \mat{1} \} $ is the set of $n\times n$ doubly stochastic matrices, $\bY \geq 0$ implies entry-wise non-negativity, $\mat{1}$ is a column vector of ones, and $\bC_\bX = \bX \bX^\top$ and $\bC_\bZ = \bZ \bZ^\top$ are $n\times n$ word covariance matrices of source and target languages, respectively. An iterative scheme is proposed for solving (\ref{eq:GW}), where each iteration involves solving an optimal transport problem with \emph{entropic regularization} \citep{peyre16a,peyre19a}. The optimal transport problem is solved with the popular \emph{Sinkhorn} algorithm \citep{cuturi13a}. 
It should be noted that the GW approach (\ref{eq:GW}) only learns $\bY$. The linear operator to map source language word embedding to the target language embedding space can then be learned by solving (\ref{eq:procrustes}). 



\noindent\textbf{Domain adaptation.} Domain adaption refers to transfer of information across domains and has been an independent research of interest in many fields including natural language processing \citep{daumeiii07a,borgwardt06a,adel17a,baktashmotlagh13a,fukumizu07a,wang15a,prettenhofer11a,wan11a,sun16a,mahadevan18a,ruder19a}.  

One modeling of interest is by \citet{sun16a}, who motivate a linear transformation on the features in source and target domains. In \citep{sun16a}, the linear map $\bA \in \R^{d\times d}$ is solved by 
\begin{equation}\label{eq:coral}
\begin{array}{ll}
\minop_{\bA \in \R^{d\times d}}	& \norm{\bA^\top \bD_\bX \bA - \bD_\bZ}_{\rm Fro}^2,
\end{array}
\end{equation}
where $\bD_1$ and $\bD_2$ are $d\times d$ are feature covariances of source and target domains (e.g., $\bD_\bX = \bX^\top \bX$ and $\bD_\bZ = \bZ^\top \bZ$), respectively. Interestingly, (\ref{eq:coral}) has a closed-form solution and shows good performance on standard benchmark domain adaptation tasks \citep{sun16a}.

\section{Domain Adaptation Based Cross-lingual Alignment}
The domain adaptation solution strategies of \citep{sun16a,mahadevan18a} can be motivated directly for the cross-lingual alignment problem by dealing with word covariances instead of feature covariances. However, the cross-lingual word alignment problem additionally has a bi-directional symmetry: if $\bY$ aligns $\bX$ to $\bZ$, then $\bY ^\top$ aligns $\bZ$ to $\bX$. 
We exploit this to propose a bi-directional domain adaptation scheme based on (\ref{eq:coral}). The key idea is to adapt the second order information of the source and the target languages into each other's domain. We formulate the above as follows: 
\begin{equation}\label{eq:proposed}
\begin{array}{lll}
\minop_{\bY \in \DS^{n}} &  \|  \bY^\top \bC_\bX \bY -  \bC_\bZ\|_{\rm Fro}^2 \\
& \quad + \ \ \|  \bY \bC_\bZ \bY^\top -  \bC_\bX\|_{\rm Fro}^2,
\end{array}
\end{equation}
%
The first term in the objective function $\|  \bY^\top \bC_\bX \bY -  \bC_\bZ\|_{\rm Fro}^2$ adapts the domain of $\bX$ (source) into $\bZ$ (target). 
Equivalently, minimizing only the first term in the objective function of (\ref{eq:proposed}) leads to row indices in $\bY^\top \bX$ aligning closely with the row indices of $\bZ$. 
Similarly, minimizing only the second term $\|  \bY \bC_\bZ \bY^\top -  \bC_\bX\|_{\rm Fro}^2$ adapts $\bZ$ (now treated as the source domain) into $\bX$ (now treated as the target domain), which means that the row indices $\bY\bZ$ and $\bX$ are closely aligned. Overall, minimizing both the terms of the objective function allows to learn the alignment matrix $\bY$ from $\bX$ to $\bZ$ and $\bY ^\top$ from $\bZ$ to $\bX$ simultaneously. Empirically, we observe that bi-directionality acts as a self regularization, leading to optimization stability and better generalization ability. 

The differences of the proposed formulation (\ref{eq:proposed}) with respect to the GW formulation (\ref{eq:GW}) are two fold. First, the formulation (\ref{eq:GW}) maximizes the inner product between $\bY^\top \bC_\bX \bY $ and $\bC_\bZ$. This inner product is sensitive to differences in the norms of $\bY^\top \bC_\bX \bY$ and $\bC_\bZ$. The proposed approach circumvents this issue since (\ref{eq:proposed})  explicitly penalizes entry-wise mismatch between $\bY^\top \bC_\bX \bY $ and $\bC_\bZ$. Second, the GW algorithm for (\ref{eq:GW}) is sensitive to choices of the entropic regularization parameter \citep{alvares18a,peyre19a}. In our case, no such regularization is required. 


Most recent works that solve optimal transport problem by optimizing over doubly stochastic matrices employ the Sinkhorn algorithm with entropic regularization \citep{cuturi13a,peyre16a,peyre19a}. 
In contrast, we exploit the Riemannian manifold structure of the set of doubly stochastic matrices ($\DS^n$) recently studied in \citep{douik19a}. $\DS^n$ is endowed with a smooth Fisher information metric (inner product) that makes the manifold smooth \citep{douik19a,sun15a,lebanon04a}. In differential geometric terms, $\DS^n$ has the structure of a Riemannian submanifold. This makes computation of optimization-related ingredients, e.g., gradient and Hessian of a function, projection operators, and retraction operator, straightforward. Leveraging the versatile Riemannian optimization framework \citep{absil08a,edelman98a,smith94a}, the constrained problem (\ref{eq:proposed}) is conceptually transformed to an \emph{unconstrained} problem over the nonlinear manifold. Consequently, most unconstrained optimization algorithms generalize well to manifolds. We solve (\ref{eq:proposed}) using the Riemannian conjugate gradient algorithm \citep{absil08a,douik19a}.

There exist several manifold optimization toolboxes such as Manopt \citep{boumal14a}, Pymanopt \citep{townsend16a}, Manopt.jl \citep{bergmann19a}, McTorch \citep{Meghwanshi_arXiv_2018} or ROPTLIB \citep{huang16a}, which have scalable off-the-shelf generic implementation of Riemannian algorithms. 
We use Manopt for our experiments, where we only need to provide the objective function (\ref{eq:proposed}) and its derivative with respect to $\bY$. The manifold optimization related ingredients are handled by Manopt internally. The computational cost per iteration of the algorithm is $O(n^2)$, which is similar to that of GW \citep{alvares18a}.

We term our algorithm as \textbf{M}anifold \textbf{B}ased \textbf{A}lignment (\algname) algorithm. Our code is available at \url{https://pratikjawanpuria.com/publications/}.

\begin{table*}[t]\centering
	{\small
		\centering
		\begin{tabular}{lrrrrrrrrrrrrr}
			\toprule
			\lhead{Method} & \thead{de-xx} & \thead{en-xx}  & \thead{es-xx} & \thead{fr-xx}  & \thead{it-xx}  & \thead{pt-xx} & \thead{xx-de} & \thead{xx-en}  & \thead{xx-es} & \thead{xx-fr}  & \thead{xx-it} & \thead{xx-pt} & \thead{avg.} \\
			\cmidrule(lr){1-1}
			\cmidrule(lr){2-7}
			\cmidrule(lr){8-13}
			\cmidrule(lr){14-14}
			GW            & 62.6 & 77.4 &\textbf{78.2} & \textbf{75.4} & \textbf{77.5}& 77.2 & 62.6 & 75.9 & \textbf{79.7} & \textbf{79.0} &\textbf{76.2} & 74.9 & 74.7 \\ 
			\textbf{\algname}    & \textbf{63.3} & \textbf{78.4}  & \textbf{78.2} & 75.3& 77.0& \textbf{77.5} & \textbf{63.1} & \textbf{77.3} & 79.4 & 78.7 & \textbf{76.2} &\textbf{75.0} & \textbf{75.0} \\
			\bottomrule
		\end{tabular}
	}
	\caption{P$@1$ for BLI on six European languages: English, German, Spanish, French, Italian, and Portuguese. Here `en-xx' refers to the average P$@1$ when English is the source language and others are target language. Similarly, `xx-en' implies English as the target language and others as source language. Thus, `avg.' shows P$@1$ averaged over all the thirty BLI results for each algorithm. The proposed algorithm {\algname} performs similar when the language pairs are closely related to each other.}\label{table:six-euro}
\end{table*}

\begin{table*}[h]\centering
	{\small
		\centering
		\begin{tabular}{lrrrrrrrrrrr}
			\toprule
			\lhead{Method} &  \thead{en-bg} & \thead{en-cs}  &  \thead{en-da} & \thead{en-el}& \thead{en-fi}&\thead{en-hu}&\thead{en-nl} & \thead{en-pl}  & \thead{en-ru} &     \\
			\cmidrule(lr){1-1}
			\cmidrule(lr){2-10}
				GW      &    22.8	 &  42.1	 &  54.4	 &  21.5 &  	37.7 &  	43.7 &  	72.9 &  	49.1 &  	36.1  &  	\\
			\textbf{\algname}  &  \textbf{38.1}  &  	\textbf{46.8}  &  	\textbf{56.1}  &  	\textbf{40.0} &  	\textbf{40.4}  &  	\textbf{46.1}  &  	\textbf{73.8}  &  	\textbf{50.4}  &  	\textbf{37.5}  &  	\\
			\midrule
			\lhead{Method} &   \thead{bg-en} & \thead{cs-en} & \thead{da-en}  & \thead{el-en} &  \thead{fi-en} &  \thead{hu-en} & \thead{nl-en} & \thead{pl-en} &  \thead{ru-en}  &  \thead{avg.}  \\
			\cmidrule(lr){1-1}
			\cmidrule(lr){2-10}
			\cmidrule(lr){11-11}
				GW     &  29.9	&52.9&	60.7&	32.7&	49.5&	57.6&	70.9&	57.7&	48.3& 47.0\\
			\textbf{\algname}  & \textbf{50.0}	& \textbf{57.7}&	\textbf{62.3}&	\textbf{54.4}&	\textbf{54.4}&	\textbf{61.0}&	\textbf{71.0}&	\textbf{60.5} & \textbf{54.1} & \textbf{53.0}\\
			\bottomrule
		\end{tabular}
	}
	\caption{P$@1$ for BLI on English and nine European languages: Bulgarian, Czech, Danish, Greek, Finnish, Hungarian, Dutch, Polish, and Russian.  The `avg.' shows P$@1$ averaged over all the eighteen BLI results. The proposed algorithm MBA outperforms GW when the bilingual mapping is learned between distant languages. }\label{table:ten_euro}
\end{table*}

\begin{table}[h]\centering
	\setlength{\tabcolsep}{5pt}
	{\small
		\centering
		\begin{tabular}{lrrrrrrr}
			\toprule
			\lhead{Method} &  \thead{en-ar} & \thead{en-hi}  &  \thead{en-tr} &  \thead{ar-en}  & \thead{hi-en} &  \thead{tr-en}  \\
			\cmidrule(lr){1-1}
			\cmidrule(lr){2-4}
			\cmidrule(lr){5-7}
			GW            &  27.4&	0.0&	40.9&	\textbf{41.0}&	0.0&	52.4\\
			\textbf{\algname}  &  \textbf{27.9} &	\textbf{25.1} &	\textbf{42.0} &	40.8 &	\textbf{28.9} &	\textbf{54.6} \\
			\bottomrule
		\end{tabular}
	}
	\caption{P$@1$ for BLI on English and three non-European languages (Arabic,  Hindi, and Turkish). MBA obtains significantly better results. 
	}\label{table:non_euro}
\end{table}

\section{Experiments}
We compare the proposed algorithm {\algname} with state-of-the-art GW alignment algorithm \citep{alvares18a} for the bilingual induction (BLI) task. Both the algorithms use second order statistics (word covariance matrices) to learn the word alignment between two languages. In our experimental setup, we first learn the word alignment between the source and the target languages and then compute cross-lingual mapping by solving the Procrustes problem (\ref{eq:procrustes}). 
For inference of nearest neighbors, we employ the cross-domain similarity local scaling (CSLS) similarity score \citep{conneau18a}. 
We report Precision$@1$ (P$@1$) as in \citep{alvares18a,artetxe18b} for the BLI task. 

We show results on the MUSE dataset \citep{conneau18a}, which consists of fastText monolingual embeddings for different languages \citep{bojanowski17} and dictionaries between several languages (but mostly with English). 
Following existing works \citep{artetxe18b,alvares18a,alaux19a}, the embeddings are normalized. 
The MUSE dataset provides predefined thirty test bilingual dictionaries between six European languages: English (en), German (de), Spanish (es), French (fr), Italian (it), and Portuguese (pt) on which we evaluate the methods. Additionally, we compute performance on the test dictionaries between English and twelve other languages: Arabic (ar), Bulgarian (bg), Czech (cs), Danish (da), Dutch (nl), Finnish (fi), Greek (el), Hindi (hi), Hungarian (hu), Polish (po), Russian (ru), and Turkish (tr). 
Following \citet{alvares18a}, we consider top $n = 20\,000$ most frequent words in the vocabulary set for all the languages during the training stage. The inference is performed on the the full vocabulary set. 

For GW, we use the original codes shared by \citet{alvares18a} and follow their recommendations on tuning the entropic regularization parameter and scaling of covariance matrices $\bC_\bX$ and $\bC_\bZ$. As a practical implementation of {\algname}, we incrementally increase $n$ starting from $1000$ to $20\,000$ every fixed-number of iterations.



We begin by discussing the results on  six close-by European languages in Table~\ref{table:six-euro}. We observe that both {\algname} and GW perform similarly when the languages are related. 
Hence, in the second set of experiments, we consider other European languages that are distant to English. 
We observe from Table \ref{table:ten_euro} that {\algname} outperforms GW, by an average BLI score of 6 points, in this challenging setting. 
Table~\ref{table:non_euro} reports results on language pairs involving English and three non-European languages. We again observe that the proposed algorithm MBA performs significantly better than GW. Overall, the experiments show the benefit of a geometric optimization framework.

\section{Conclusion}
Aligning the metric spaces of languages has a wide usage in cross-lingual applications. A popular approach in literature is the Gromov-Wasserstein (GW) alignment approach \citep{memoli11a,peyre16a,alvares18a}, which constructs a transport map by viewing the two embedding spaces as distributions. {In contrast, we have viewed} unsupervised bilingual word alignment as an instance of the more general unsupervised domain adaptation problem. {In particular, our formulation allows search over the space of doubly stochastic matrices and induces bi-directional mapping between the source and target words. Both are motivated solely from the language perspective. The Riemannian framework allows to exploit the geometry of the doubly stochastic manifold.} Empirically, we observe that the proposed algorithm {\algname} outperforms the GW algorithm for learning bilingual mapping \citep{alvares18a}, demonstrating the benefit of geometric optimization modeling. 




\bibliography{../multilingual-unsupervised-draft/main}
\bibliographystyle{acl_natbib}

\end{document}